### v1

# A time-to-digital converter with steady calibration through single-photon detection

Abstract
Time-to-Digital Converters (TDCs) are a crucial tool in a wide array of fields, in particular for quantum communication, where time taggers performance can severely affect the quality of the entire application. Nowadays, FPGA-based TDCs present a viable alternative to ASIC ones, once the nonlinear behaviour due to the intrinsic nature of the device is properly mitigated. To compensate said nonlinearities, a calibration procedure is required, usually based on an interpolation methods. Here we present the design and the demonstration of a TDC that is FPGA-based and showing a residual jitter of 27 ps, that is scalable for multichannel operation. The application in Quantum Key Distribution (QKD) is discussed with a unique calibration method based on the exploitation of single-photon detection that does not require to stop the data acquisition or to use any interpolation methods, thus increasing accuracy and removing data loss. The calibration was tested in a relevant environment, investigating the device behaviour between 5°C and 80°C. Moreover, our design is capable of continuously streaming up to 12 Mevents/s for up to $\sim$1 week without the TDC overflowing.
I Introduction
Time-to-digital converters (TDCs) are used in a wide range of applications [1, 2] where picosecond resolution is needed for time detection, such as high-energy particle physics [3], light detection and ranging (LiDAR) [4], quantum communications [5, 6, 7] and quantum random number generation [8]. Recently, Field-Programmable Gate Arrays (FPGAs) have emerged as a popular platform for implementing TDCs due to their flexibility, high-speed performance, and cost-effectiveness [9, 10, 11, 12]. However, standard FPGA-based TDCs require periodic calibration to maintain accuracy, which requires either interpolation methods or stopping the data acquisition process, leading to potential data loss and reduced efficiency [13]. Although it is possible to optimize and automatize the calibration procedure [14, 15, 16], to the best of our knowledge, there are no proposals to overcome the data losses.
Quantum Key Distribution (QKD) allow two parties (Alice and Bob) to create a shared, secret key. To do so, Alice encodes the information in degrees of freedom of a given quantum system, usually photons [17]. To make the sharing of the key successful, Bob requires precise timing of arrival of photons to synchronize both parties effectively, and to allow better distinguishability between signal and noise. For example, algorithms like Qubit4Sync allow Bob to reconstruct Alice’s clock signal from his own qubit measurements [18]. Moreover, QKD is currently evolving to be implemented using satellite links to increase the distance achievable with this protocols [19]. This implies that communication could only be done while the satellite passes through the ground station’s field of view. During this limited time, Bob should receive as much data as possible while maintaining its performance, implying that both calibrating or losing calibration mid-measurement can considerably hinder key generation.
In this work, we propose an FPGA-based TDC, named MARTY and implemented on a Zynq7020, that addresses the calibration challenge by integrating a so-called steady calibration [^1] We chose the term “steady calibration” as the terms “self-calibration”, “automatic calibration” or “real-time calibration” are standard forms to indicate just an online calibration procedure which usually include a phase of data loss. method, implemented by exploiting single-photon detection, similarly to what is done in [20] but in a different application framework. This method allows the TDC to calibrate itself using the data it acquires, eliminating the need to stop the acquisition process. This feature not only enhances the efficiency of the system, but also improves the overall time accuracy as each measurement is also a contribution to calibration and there is no distinction between “data-acquisition” and “calibration” phases.
The paper has the following structure: in Section II we present the working principle of the system, including the FPGA design, normal and steady calibration procedures, and finally how we performed temperature variation tests on the device; in section III we present the results obtained in this work, including a QKD test performed with MARTY; eventually, in section IV we present the final conclusions and future perspectives.
II Materials and methods
Figure 1: Functioning principle of a steadily-calibrated TDC (right) compared with a normal TDC (left).
II.1 Design of the device
A generic TDC channel can consists of two components: a clock counting module for the coarse time estimation and a tapped delay line (TDL) for the fine estimation (Fig. 1 left).
While the clock counting module is a binary counter that increases at each rising edge of a clock, the TDL is a “propagation line” where the electric signal travels through a chain of similar elements, each one with an intrinsic propagation delay. The implementation of such structure can be realized in different ways. In MARTY, for the coarse counting module, the DSP48 components were used as they grant high efficiency and speed, clocked by a $f_{\rm s}=412.5$ MHz clock, $\mathrm{\Phi}_{\mathrm{sample}}$. With the given clock frequency and the choice of encoding the coarse counter in 48 bits, MARTY has a total runtime of $2^{48}/f_{\rm s}\sim 1$ week before overflow.
The TDL was implemented through a chain of fast-carry components, the CARRY4 element (each containing 4 fast-carry). As usual in TDC implementation, each point of the TDL is flip-flopped in order to snapshot the current state at each clock edge. This snapshot output can be considered as a thermometer code that characterizes the propagation time inside the TDL and can be translated into a binary code using a dedicated decoder. The decoding algorithm used in MARTY was the one presented by Adamič et al. in [21], which is based on a pipelined adder tree. This algorithm was chosen due to its high robustness against ‘bubble errors’ without increasing the complexity of the system. ‘Bubbles’ behave like a ‘bit swap’ between two neighbouring bits of the thermometer code (e.g. a code that should be 11111000 instead appears as 11110100), and they appear due to a variety of effects such as metastability on the flip-flops used to snapshot the TDL state [1].
It is worth to notice that the propagation time inside each fast-carry component is not uniform, ranging from units of picoseconds to around $100$ ps. Therefore, a calibration is needed to precisely characterize such propagation times. By using 36 CARRY4 elements (144 fast-carry components in total), the coarse period $\tau_{\rm coarse}=1/f_{s}\simeq 2.42$ ns was fully covered. Indeed, a maximum length of $N_{c}\in[130,136]$ bins was obtained in all data acquisitions.
The final timestamp is obtained by combining the binary value of the TDL thermometer code, named fine value, and the value obtained from the clock counting module, named coarse value. The output timestamps of a channel can then be stored in a block RAM (BRAM) module to later transfer them to an external PC.
The device takes advantage of the System-on-a-Chip (SoC) technology, exploiting not only the FPGA but also the CPU(s) counterpart to implement a real-time data transfer from the BRAM to the PC via Ethernet protocol [22]. This scheme allows to continuously stream the TDC data from the FPGA to the PC with the only limitation of the Ethernet capability, bringing to an acquisition speed of up to 15 Mevents/s without data loss. Furthermore, to facilitate software integration, a request-based scheme was also implemented, as it is a standard approach on commercial TDCs to transfer data to a PC. In this scheme, the acquired data is sent only when receiving an external request from a PC.
Figure 2: (a) Histogram obtained with a code density test using an FPGA-based RO as input. (b) Percentile difference of the histogram obtained with a code density test using a laser detection signal with respect to the RO one.
II.2 System calibration and auto-calibration
The calibration method is based on the evaluation of the results of a code density test. A code density test consists of using a signal uniformly distributed over the time spanned by the TDL ($1/f_{s}\approx 2.42$ ns), and measuring the number of occurrences obtained from each bin. If the input signal is uniformly distributed and enough events were recorded, the occurrences in each bin should be proportional to its propagation time [2]. As a standard solution, the code density test is performed using an internal Ring Oscillator (RO) [^2] A ring oscillator consists of an unstable loop of delay elements (usually not-gate) that generates a random output., that generates a uniformly-distributed signal connected to the input of the TDC. To obtain the minimum number of events needed to calibrate, $N_{events}$, the method proposed in [23] was used, such that
$N_{events}>\left(\frac{z_{\alpha/2}}{\beta}\right)^{2}\cdot(M-1),$ (1)
where $z_{\alpha/2}$ corresponds to the area under the standard normal distribution curve of $\alpha/2$, $M=8$ the number of bits, $\beta=0.02$ represents a tolerance of 2% and $\alpha=0.01$ represents a confidence interval of 99%, from which the required number of events $N_{events}>116110$ was obtained.
From the number of events detected in the bin $i$, $w_{i}$, it is possible to derive the propagation time of the calibrated bin $\delta t_{i}=\frac{w_{i}}{\sum_{j}w_{j}}\tau_{\rm sample}$, with $i\in[1,N_{c}]$ (Fig. 2). Then, the calibrated time is defined as
$t^{(\rm c)}_{i}=\frac{1}{2}\delta t_{i}+\sum_{j<i}\delta t_{j}.$ (2)
This type of calibration requires the user to stop the current acquisition for around 30 ms, switching from the input signal to the RO one.
In this work, we propose an extension of this method, called steady calibration, which allows the system to continuously calibrate while the device is acquiring data (Fig. 1 right). This method consists of storing the code density histogram in an ordered manner, so that at each new detected event, the first one of the histogram is deleted and the new one is appended at the end. Then, the calibration curve is recalculated with the new data.
The steady calibration method allows the exploitation of the acquired TDC data to calibrate itself in real time, assuming that the detected events are uniformly distributed over the TDL. Considering a pulsed laser as an event source with the repetition rate $f_{q}$ disciplined by a clock $\mathrm{\Phi}_{\mathrm{qubit}}$, the temporal distribution is expected not to be uniform in the TDL outputs only if the two clocks $\mathrm{\Phi}_{\mathrm{qubit}}$ and $\mathrm{\Phi}_{\mathrm{sample}}$ are commensurable, or locked to each other, and perfectly stable without any drift or jitter. Therefore, the combination of non commensurable clock frequencies and drift effects, can assure a uniform distribution of the input events.
To test the proposed method, we obtained the calibration curve using a laser source and a single photon detector (SPD) in place of the integrated RO. When comparing both histograms obtained with a code density test (Fig. 2), the difference between both is considered negligible, with a $\chi^{2}=0.026$. From this, we can conclude that using an external single photon detector with a laser signal is equivalent to the integrated RO in the FPGA chip for calibration purposes.
II.3 Temperature test setup
The validity of the steady calibration method was verified by forcing a change in the working point of the system, which should require a calibration. Therefore, the device was tested at different temperatures as this can have a relevant impact on the performance of a TDC and is considered strictly related to the TDC calibration process [24]. To modify the device temperature, we used a combination of a Peltier cell and an RAL 9006 climatic chamber by Angelatoni Test Technologies. The combination of both temperature control systems allowed the FPGA chip to reach a broader temperature range, from 5 °C to 80 °C. To measure the temperature of the chip, the included XADC sensor of the board was used, which gives a temperature measurement in the center of the chip itself.
The jitter and resolution of a two-channel implementation of MARTY was used to characterize its performance. To calculate the jitter, an input signal is divided into the two channels, which we assume equal for this purpose, and the histogram of the differences between the timetags is obtained, from which the jitter can be obtained. For this approximation, we assume that both channels have the same jitter and that the jitter of the source does not affect the resulting histogram due to both channels receiving the same input signal.
To perform the temperature measurements, a 1550 nm attenuated laser source was used as an input to a single-photon detector (PMD-IR detector by MPD), which was then divided and used as input for both channels of the TDC. The chip temperature was scanned in the whole range of 5 °C to 80 °C with a step of 1 °C, where for each temperature the jitter and resolution were evaluated.
II.4 Quantum Key Distribution test setup
To evaluate the device performance for QKD applications, while also comparing it with commercial TDCs, a QKD transmitter like the one proposed in [25] was used to encode an “HVDD” sequence. [^3] An “HVDD” sequence is a sequence of four qubits encoded using single-photon with Horizontal-Vertical-Diagonal-Diagonal polarizations. We used a standard polarization receiver scheme with four PMD-IR detectors by MPD, where for the time-tagger we used both a QuTAG from QuTools and MARTY. To evaluate the performance, a detection peak was measured with both devices, and the Quantum Bit Error Rate (QBER) for the HVDD sequence was estimated.
III Results and discussions
Figure 3: DNL (in blue) and $\text{DNL}^{t}$ (in red) obtained for a single channel of the TDC at room temperature.
Figure 4: Calibrated time as a function of the bin number of the TDC, in orange the ideal behaviour for a uniform delay line.
MARTY’s main performance metrics, such as resolution, differential nonlinearity (DNL), and jitter were characterized using a two-channel TDC configuration. The DNL represents the deviation of the delay line with respect to an ideal one (uniformly distributed in time) and it was evaluated through a code density test with an internal RO as the input of the TDC, obtaining the counts for bin $i$, $w_{i}$ (Fig. 2a), from which the DNL was obtained as
$\text{DNL}_{i}=\frac{w_{i}-\langle w\rangle}{\langle w\rangle},$ (3)
[…]
$\tau_{\mathrm{res}}=\frac{t^{(c)}_{N_{c}}}{N_{c}}=18.22~{}\text{ps}$ (4)
was obtained at room temperature. Moreover, once the bins were properly calibrated, the DNL was evaluated again, this time using calibrated times $t^{(c)}$ as
${\rm DNL}^{t}_{i}=\begin{cases}0&\quad\text{if i=1},\\ \frac{t^{(c)}_{i}-t^{(c)}_{i-1}-\langle\Delta t^{(c)}\rangle}{\langle\Delta t^{(c)}\rangle}&\quad\text{if i}>1,\\ \end{cases}$ (5)
where $\langle\Delta t^{(c)}\rangle=\langle t_{i+1}^{(c)}-t_{i}^{(c)}\rangle=\tau_{\mathrm{res}}$.
Lastly, to characterize the jitter of the system, a scheme like the one described in section II.3 was used, paired with a square wave as an input for both channels at room temperature. To obtain the jitter, a Gaussian fit was applied to the histogram of the difference between consecutive tags, from which the jitter was estimated as the $\text{FWHM}=27.11$ ps (Fig. 5). We notice that the jitter is obtained by dividing by $\sqrt{2}$ the standard deviation of the Gaussian fit.
Figure 5: Histogram of the difference between two consecutive timetags in a two-channel implementation. In red, the corresponding Gaussian fit, from which the jitter was estimated.
To determine the device performance with respect to environment changes, a temperature characterization of the system using the setup described in Section II.3 was performed. The data acquisition process consisted on storing, for each temperature, a file with the calibration curve for each channel along with the registered tags in non-calibrated format, with their respective channel. This format allowed to analyze the data with or without calibration, while also letting the possibility to implement calibration at particular points in the measurement. The acquired data for each temperature was processed in four different ways:
  1. 1.
    Calibrated using the calibration curve obtained with the RO at 5 °C (light orange curve in Fig. 7).
  2. 2.
    Calibrated using the first time tags registered with the single photon detector at 5 °C (light blue curve in Fig. 7).
  3. 3.
    Calibrated using the calibration curve obtained for each temperature (red curve in Fig. 7).
  4. 4.
    Steadily calibrated (blue curve in Fig. 7).
Figure 6: Mean single-bin propagation time for each channel of the TDC as a function of the temperature of the chip. The different offset between both channels was attributed to non-uniformity of components in the chip.
Figure 7: Comparison of the jitter of the system as a function of temperature obtained for different calibration methods: with the RO at 5 °C and maintained through the scan (light orange); with the RO every 1 °C step (red); with the single photon detector at 5 °C and maintained through the scan (light blue); and steadily with the single photon detector (blue).
As a first step, the relation between temperature and the average propagation time inside the delay line, $\langle\delta t\rangle$, was analysed. Since the maximum number of registered bins is temperature and channel dependent (obtaining around 130 bins for 5 °C and up to 136 for 80 °C), we restricted this analysis to the first 120 elements of the delay line, such that
$\langle\delta t\rangle=\frac{\sum_{j=1}^{120}\delta t_{j}}{120}.$ (6)
The observed behavior depends on the particular architecture and materials of the Zynq7020 chip (28 nm process), of which not much information is publicly available, concluding that the CARRY4 element has a propagation time that is, in average, inversely proportional to temperature.
Then, the jitter was calculated in the 4 scenarios previously mentioned (Fig. 7). For both cases in which the calibration was done just at 5 °C, increasing the chip temperature monotonically increased the jitter of the system (as also verified by Alshehry et al. in [26]), which highlights the importance of the steady calibration scheme, since the transistors’ characteristics are temperature dependent. On the other hand, when applying a calibration for every temperature step, the jitter decreased until 30 °C, which implies an improvement in performance, and then increased until it stabilized at 60 °C, close to the starting value. Finally, the one obtained with the steady calibration had a very similar shape to the one obtained by calibrating at each temperature, but with two notable improvements. In first instance, the system’s jitter was found to be more stable, presenting an average standard deviation $\langle\sigma_{\text{steady}}\rangle=0.64$ ps, compared to the $\langle\sigma_{\text{RO}}\rangle=1.33$ ps obtained for the case when RO calibration was performed before each measurement. Moreover, as previously mentioned, the steady calibration does not require the acquisition to be stopped at any time, thus allowing the system to not lose any data.
Finally, a test with a QKD polarization-encoded source and receiver scheme was performed. The width of the detected pulse depends on the width of the optical pulse itself, the jitter of the detector, and the jitter of the time tagger. For both devices, the mean QBER obtained was of $2.2$ [^4] The QKD source and receivers were not properly optimized for performance as it is outside the scope of this work., showing that our design presents good performance for QKD applications. We compared the width of the detected pulse using both MARTY and the QuTAG, a high-level commercial time tagging system with 13.4 ps jitter on the standard version (Fig. 8). It was observed that, even though Marty’s jitter was higher than the QuTAG, the pulse width obtained with both devices were comparable. This behaviour is due to most of the width of the pulse coming from the detector’s jitter, which is considerable higher than the time-tagger ($\sim 100$ ps). Indeed, moving to a higher level FPGA technology (e.g., Ultrascale+ with 16 nm FinFET process) can considerably increase the device’s performances [27].
Figure 8: Pulse shape obtained with MARTY (orange bars) and the QuTAG (blue bars), with their corresponding Gaussian fit.
IV Conclusion
[…]
Acknowledgment
M.R.B.W. acknowledges support from the European Union’s Horizon Europe Framework Programme under the Marie Sklodowska Curie Grant No. 101072637, Project Quantum-Safe Internet (QSI).
M.R.B.W. acknowledges funding from the italian government via a study scholarship granted by the Ministry of Foreign Affairs and International Cooperation (MAECI) A.Y. 2021-2022.
This work is (partially) supported by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU; Agenzia Spaziale Italiana, project Q-SecGroundSpace (Accordo n. 2018- 14-HH.0, CUP: E16J16001490001); Agenzia Spaziale Italiana (2020-19- HH.0 CUP Grant No. F92F20000000005, Italian Quantum CyberSecurity I-QKD).
The authors would like to thank Dr. S. Bonora for the usage of the climatic chamber and Mr. A. Vanzo for the help in setting it up.